OpenStudy (anonymous):

A student is 5 feet, 2 inches tall. What is her height in meters?

OpenStudy (anonymous):

62 inches or 5' 2 " = 1.5748 meters

OpenStudy (anonymous):

can you show me the work? :)

OpenStudy (anonymous):

like how did you convert it?

OpenStudy (anonymous):

1 Foot = 12 inches So, 5 Feet=5*12=60 inches. Total height= (60+2) inches (since he is 5 feet 2 inches tall) Height=62 inches. Now, 1 Meter = 39.3700787 Inches So, 1 inch=1/39.37 m So, 62 inches= 62/39.37 m = 1.57 m Thats your answer :)